The perimeter of a rectangle is 54 meters.The longer side is three more than the shorter side.Find the length of each side.

Answer by mananth(16949)

let length = l
width == w
l=w+3
2(l+w)=54
substitute l
2((w+(w+3))= 54
2(2w+3)=54
4w+6=54
4w=48
w= 12
l= 12+3=15
check
2(15+12)
2*27
54